### v2

# Two-pion exchange nucleon-nucleon potentials with Roper resonance excitation

Abstract
Motivated by the recent paper arXiv:2602.11815 [nucl-th], we calculate in these notes the spectral functions (i.e. imaginary parts) of the NN-potentials as they arise from $2\pi$-exchange with single and double Roper resonance excitation. In contrast to the full one-loop calculation, the spectral functions of the isoscalar and isovector central and tensor potentials can be given in simple analytical form. The pertinent momentum-space potentials $V_{C}(q),W_{C}(q),V_{T}(q),W_{T}(q)$ are obtained via subtracted dispersion relations. This representation allows also to include a regulator function, that tames high-momentum components of the chiral $2\pi$-exchange. The calculation is extended to $2\pi$-exchange with combined $\Delta(1232)$-isobar and Roper resonance excitation.
1 Preparation
In the nonrelativistic treatment the nucleon-nucleon interaction as it arises from two-pion exchange takes the following form:
$T_{N\!N}=V_{C}(q)+\vec{\tau}_{1}\!\cdot\!\vec{\tau}_{2}\,W_{C}(q)-\big[V_{T}(q)+\vec{\tau}_{1}\!\cdot\!\vec{\tau}_{2}\,W_{T}(q)\Big](\vec{\sigma}_{1}\!\times\!\vec{q}\,)\!\cdot\!(\vec{\sigma}_{2}\!\times\!\vec{q}\,)\,$, (1)
where $\vec{\tau}_{1,2}$ and $\vec{\sigma}_{1,2}$ denote the isospin- and spin-operators of both nucleons and $\vec{q}$ is the momentum transfer with magnitude $q=|\vec{q}\,|$. The sign-convention for $T_{N\!N}$ has been chosen such that one-pion exchange gives $T_{N\!N}^{(1\pi)}=(g_{A}/2f_{\pi})^{2}\vec{\tau}_{1}\!\cdot\!\vec{\tau}_{2}\,\vec{\sigma}_{1}\!\cdot\!\vec{q}\,\vec{\sigma}_{2}\!\cdot\!\vec{q}\,(m_{\pi}^{2}+q^{2})^{-1}$, with $g_{A}\simeq 1.3$, $f_{\pi}=92$ MeV (pion decay constant), and $m_{\pi}=138\,$MeV the average pion mass. The long-range parts of the $2\pi$-exchange NN-potentials are obtained from the corresponding spectral functions (i.e. imaginary parts) via twice- or once-subtracted dispersion relations as:
$V_{C}(q)=C_{0}+C_{1}\,q^{2}+{2q^{4}\over\pi}\!\int_{2m_{\pi}}^{\infty}\!\!d\mu\,{{\rm Im}V_{C}(i\mu)\over\mu^{3}(\mu^{2}+q^{2})}\,,\qquad V_{T}(q)=C_{T}-{2q^{2}\over\pi}\!\int_{2m_{\pi}}^{\infty}\!\!d\mu\,{{\rm Im}V_{T}(i\mu)\over\mu(\mu^{2}+q^{2})}\,.$ (2)
For the isovector central and tensor potentials $W_{C}(q),W_{T}(q)$ analogous dispersion relations hold. As a further modification, that is often used, this representation allows to introduce a regulator function, e.g. $f_{\Lambda}(\mu)=\exp(-\mu^{2}/\Lambda^{2})$, in the spectral integral to tame high-momentum components of the chiral $2\pi$-exchange interaction.
The imaginary part of a one-loop $2\pi$-exchange diagram (see Fig. 1) is calculated with the help of the Cutkosky cutting rule as [1]:
${\rm Im}\int\!{d^{4}l\over(2\pi)^{4}i}\,{A\otimes B\over(m_{\pi}^{2}-l\!\cdot\!l)(m_{\pi}^{2}-(q\!-\!l)\!\cdot\!(q\!-\!l))}={k\over 16\pi\mu}\int_{-1}^{1}\!dx\,A\otimes B\,$, (3)
where AA and $B$ symbolize vertices on the left and right hand side of the diagram and the angular integral $\int_{-1}^{1}\!dx$ effectively goes over the Lorentz-invariant $2\pi$-phase space. Furthermore, $\mu=\sqrt{q\!\cdot\!q}\geq 2m_{\pi}$ denotes the $2\pi$-invariant mass and $k=\sqrt{\mu^{2}/4-m_{\pi}^{2}}$ is the (single) pion center-of-mass momentum. For the $\bar{N}\!N\to\pi\pi\to\bar{N}\!N$ $t$-channel kinematics under consideration in the heavy baryon limit, one sets the nucleon velocity four-vector $v^{\nu}$ equal to $v^{\nu}=(0,i\,\vec{v}\,)$ with $\vec{v}$ a 3-component unit-vector, which leads to the relations $v\!\cdot\!q=0$ and $v\!\cdot\!l=-i\,x\,k$. In the first step the emerging integrand is complex-valued, but it becomes real-valued after symmetrizing it with respect to $x\to-x$, and leaves an integral $\int_{0}^{1}\!dx$ over positive values of the cosine $x$. The intermediate state nucleon is handled with an infinitesimally small positive mass-splitting $\eta$, which leads to $\lim_{\eta\to 0}\eta/(\eta^{2}+k^{2}x^{2})=\pi\delta(x)/k$ and $\int_{0}^{1}\!dx\,\delta(x)=1/2$. Further technical details on the method can be found in section 2.2 of ref.[1], but note that therein the numerator of eq.(9) must be corrected to $\sqrt{\mu^{2}-4m_{\pi}^{2}}$.
The $\pi NN^{*}$ vertex and heavy Roper propapator pertinent to the present calculation read [2]:
${g^{\prime}_{A}\over 2f_{\pi}}\,\vec{\sigma}\!\cdot\!\vec{l}\,\,\tau^{a}\,,\quad{i\over v\!\cdot\!l-\rho}\,,\qquad\qquad{3g_{A}\over 2\sqrt{2}f_{\pi}}\,\vec{S}\!\cdot\!\vec{l}\,\,T^{a}\,,\quad{i\over v\!\cdot\!l-\Delta}\,$, (4)
with $g^{\prime}_{A}$ a coupling constant (various values are discussed in section III.A of ref.[2]) and $\rho\simeq 500\,$MeV the mass difference between the Roper $N^{*}(1440)$-resonance and the nucleon. The right half of eq.(4) gives the analogous input for $\Delta(1232)$-isobar excitation [3], where the $\Delta N$-mass difference is smaller: $\Delta\simeq 300\,$MeV. The $2\times 4$ spin- and isospin-transition matrices $S_{i}$ and $T_{a}$ fulfill the relations $3S_{i}S_{j}^{\dagger}=2\delta_{ij}-i\epsilon_{ijk}\sigma_{k}$ and $3T_{a}T_{b}^{\dagger}=2\delta_{ab}-i\epsilon_{abc}\tau_{c}$. The coupling ratio $3/\sqrt{2}$ to $\pi N\!N$ is empirically well satisfied.
2 Spectral functions of $2\pi$-exchange NN-potentials
In this section we present the expressions for the spectral functions $V_{C}(i\mu), Im$$W_{C}(i\mu)$, $V_{T}(i\mu), Im$$W_{T}(i\mu)$ as they result from the one-loop $2\pi$-exchange diagrams shown in Fig. 1. We remind of the abbreviation $k$ with its meaning: $2k=\sqrt{\mu^{2}-4m_{\pi}^{2}}$.
Figure 1: One-loop $2\pi$-exchange diagrams with single and double Roper resonance excitation.
Triangle diagrams with Roper excitation
The (left) triangle diagram in Fig. 1 and its mirror partner lead to the following (doubled) contribution to the imaginary part of the isovector central potential:
${\rm Im}W_{C}(i\mu)={g^{\prime 2}_{A}\over\pi\mu(2f_{\pi})^{4}}\bigg\{\Big({2m_{\pi}^{2}\over 3}-{5\mu^{2}\over 12}-\rho^{2}\Big)k+\rho\Big({\mu^{2}\over 2}-m_{\pi}^{2}+\rho^{2}\Big)\arctan{k\over\rho}\bigg\}\,.$ (5)
Since this diagram involves on one side the isovector Weinberg-Tomozawa $2\pi N\!N$-contact vertex, there is no $g_{A}^{2}$ in the prefactor, as it was incorrectly written in eq.(A1) of ref.[2].
Planar and crossed box diagrams with single Roper excitation
The (central) planar and crossed box diagrams in Fig. 1 with single Roper excitation and their mirror partners lead to the following (doubled) contributions to the spectral functions:
${\rm Im}V_{C}(i\mu)={3g_{A}^{2}g^{\prime 2}_{A}\over\mu(4f_{\pi})^{4}\rho}(\mu^{2}-2m_{\pi}^{2})^{2}\,,\qquad V_{C}(q)={6g_{A}^{2}g^{\prime 2}_{A}\over\pi(4f_{\pi})^{4}\rho\,q}(2m_{\pi}^{2}+q^{2})^{2}\arctan{q\over 2m_{\pi}}\,$, (6)
${\rm Im}W_{C}(i\mu)={g_{A}^{2}g^{\prime 2}_{A}\over\pi\mu(2f_{\pi})^{4}}\bigg\{\Big({11\mu^{2}\over 12}-{5m_{\pi}^{2}\over 3}+\rho^{2}\Big)k-{1\over\rho}\Big({\mu^{2}\over 2}-m_{\pi}^{2}+\rho^{2}\Big)^{2}\arctan{k\over\rho}\bigg\}\,$, (7)
${\rm Im}V_{T}(i\mu)={3g_{A}^{2}g^{\prime 2}_{A}\over\pi\mu(4f_{\pi})^{4}}\bigg\{{1\over\rho}(\mu^{2}-4m_{\pi}^{2}+4\rho^{2})\arctan{k\over\rho}-4k\bigg\}\,$, (8)
${\rm Im}W_{T}(i\mu)={g_{A}^{2}g^{\prime 2}_{A}\over\mu(4f_{\pi})^{4}\rho}(4m_{\pi}^{2}-\mu^{2})\,,\qquad W_{T}(q)={2g_{A}^{2}g^{\prime 2}_{A}\over\pi(4f_{\pi})^{4}\rho\,q}(4m_{\pi}^{2}+q^{2})\arctan{q\over 2m_{\pi}}\,.$ (9)
Modulo additive constants (plus const $q^{2}$) and the opposite sign-convention, the results for $V_{C}(q)$ and $W_{T}(q)$ agree with the expressions written in eqs.(5,14) of ref.[2].
Planar and crossed box diagrams with double Roper excitation
The (right) planar and crossed box diagrams in Fig. 1 with double Roper excitation lead to the following contributions to the spectral functions:
${\rm Im}V_{C}(i\mu)$ $=$ ${3g^{\prime 4}_{A}\over\pi\mu(4f_{\pi})^{4}}\bigg\{\bigg[2\mu^{2}+4\rho^{2}+{8(\rho^{2}-m_{\pi}^{2})^{2}\over\mu^{2}-4m_{\pi}^{2}+4\rho^{2}}\bigg]k$ (10)
$+{1\over\rho}\Big({\mu^{2}\over 2}-m_{\pi}^{2}+\rho^{2}\Big)(\mu^{2}-2m_{\pi}^{2}-6\rho^{2})\arctan{k\over\rho}\bigg\}\,$,
${\rm Im}W_{C}(i\mu)$ $=$ ${g^{\prime 4}_{A}\over\pi\mu(4f_{\pi})^{4}}\bigg\{\bigg[{2\over 3}(17\mu^{2}-20m_{\pi}^{2})+16\rho^{2}+{16(\rho^{2}-m_{\pi}^{2})^{2}\over\mu^{2}-4m_{\pi}^{2}+4\rho^{2}}\bigg]k$ (11)
$+{1\over\rho}(2m_{\pi}^{2}-2\rho^{2}-\mu^{2})(\mu^{2}-2m_{\pi}^{2}+10\rho^{2})\arctan{k\over\rho}\bigg\}\,$,
${\rm Im}V_{T}(i\mu)={3g^{\prime 4}_{A}\over\pi\mu(4f_{\pi})^{4}}\bigg\{{1\over\rho}\Big({\mu^{2}\over 4}-m_{\pi}^{2}+3\rho^{2}\Big)\arctan{k\over\rho}-3k\bigg\}\,$, (12)
${\rm Im}W_{T}(i\mu)={g^{\prime 4}_{A}\over\pi\mu(4f_{\pi})^{4}}\bigg\{{1\over\rho}\Big(2m_{\pi}^{2}+2\rho^{2}-{\mu^{2}\over 2}\Big)\arctan{k\over\rho}-2k\bigg\}\,.$ (13)
The spectral functions for the $2\pi$-exchange NN-potentials with single and double $\Delta(1232)$-isobar excitation (computed in section 3 of ref.[3]) are obtained from the expressions written here in eqs.(5-13) by replacing the coupling constant $g^{\prime}_{A}\to g_{A}$, the mass-splitting $\rho\to\Delta$, and multiplying with weight factors 1,2,1,1,1/2,4,1,1,1/4-1,2,-1,-1,1/2,4,1,1,1/4, in that order.
Planar and crossed box diagrams with combined Roper and Delta excitation
As emphasized in ref.[2], when including the Roper $N^{*}(1440)$-resonance one must also consider explicit contributions from $\Delta(1232)$-isobar excitation because of the lower $\Delta N$-mass splitting. The NN-potentials as they arise from single and double $\Delta(1232)$-excitation have been given in section  3 of ref.[3]. Here, we complete the pattern of $2\pi$-exchange with intermediate low-lying nucleon-resonance excitations by considering the combination of the $\Delta(1232)$-isobar and the Roper resonance. The (right) planar and crossed box diagrams in Fig. 1 with combined $\Delta(1232)$-isobar and Roper excitation and their mirror partners lead to the following (doubled) contributions to the spectral functions:
${\rm Im}V_{C}(i\mu)$ $=$ ${3g_{A}^{2}g^{\prime 2}_{A}\over 64\pi\mu f_{\pi}^{4}}\bigg\{4\rho\Delta\,k+{\rho\over\rho^{2}\!-\!\Delta^{2}}(\mu^{2}-2m_{\pi}^{2}+2\Delta^{2})^{2}\arctan{k\over\Delta}$ (14)
$-{\Delta\over\rho^{2}\!-\!\Delta^{2}}(\mu^{2}-2m_{\pi}^{2}+2\rho^{2})^{2}\arctan{k\over\rho}\bigg\}\,$,
${\rm Im}W_{C}(i\mu)$ $=$ ${g_{A}^{2}g^{\prime 2}_{A}\over 64\pi\mu f_{\pi}^{4}}\bigg\{\Big[{1\over 3}(20m_{\pi}^{2}-11\mu^{2})-4\Delta^{2}-4\rho^{2}\Big]k$
$-{\Delta\over\rho^{2}\!-\!\Delta^{2}}(\mu^{2}-2m_{\pi}^{2}+2\Delta^{2})^{2}\arctan{k\over\Delta}+{\rho\over\rho^{2}\!-\!\Delta^{2}}(\mu^{2}-2m_{\pi}^{2}+2\rho^{2})^{2}\arctan{k\over\rho}\bigg\}\,$,
${\rm Im}V_{T}(i\mu)$ $=$ ${3g_{A}^{2}g^{\prime 2}_{A}\over\pi\mu(4f_{\pi})^{4}}\bigg\{4k+{\Delta\over\rho^{2}\!-\!\Delta^{2}}(\mu^{2}-4m_{\pi}^{2}+4\Delta^{2})\arctan{k\over\Delta}$ (16)
$-{\rho\over\rho^{2}\!-\!\Delta^{2}}(\mu^{2}-4m_{\pi}^{2}+4\rho^{2})\arctan{k\over\rho}\bigg\}\,$,
${\rm Im}W_{T}(i\mu)={g_{A}^{2}g^{\prime 2}_{A}\over\pi\mu(4f_{\pi})^{4}}\bigg\{{\Delta\over\rho^{2}\!-\!\Delta^{2}}(\mu^{2}-4m_{\pi}^{2}+4\rho^{2})\arctan{k\over\rho}-{\rho\over\rho^{2}\!-\!\Delta^{2}}(\mu^{2}-4m_{\pi}^{2}+4\Delta^{2})\arctan{k\over\Delta}\bigg\}\,$, (17)
which are obviously symmetric under the permutation of the mass-splittings $\Delta\leftrightarrow\rho$. The apparent complexity of one-loop integrals with heavy baryon propagators involving different mass-splittings boils down on the level of the spectral function to symmetrizing a term involving $\Delta^{-1}\!\arctan(k/\Delta)$ by another term involving $\rho^{-1}\!\arctan(k/\rho)$.
By comparison with the analytical expressions given in section 3 of ref.[3], one learns that the translation of $\mu$-dependent spectral functions into $q$-dependent $2\pi$-exchange NN-potentials proceeds as follows:
$\mu^{2}\to-q^{2}\,,\qquad\quad{k\over\mu}\to-{L(q)\over\pi}\,,\quad L(q)={\sqrt{4m_{\pi}^{2}+q^{2}}\over q}\ln{\sqrt{4m_{\pi}^{2}+q^{2}}+q\over 2m_{\pi}}\,$,
${k\over\mu(\mu^{2}-4m_{\pi}^{2}+4\rho^{2})}\to{L(q)-L(2\sqrt{\rho^{2}-m_{\pi}^{2}}\,)\over\pi(4m_{\pi}^{2}+q^{2}-4\rho^{2})}\,,\qquad\quad{1\over\mu\,\Delta}\arctan{k\over\Delta}\to{2\over\pi}D_{\Delta}(q)\,$, (18)
and an analogous function $D_{\rho}(q)$ [3], if the mass-splitting $\Delta\simeq 300\,$MeV is replaced by $\rho\simeq 500\,$MeV.